# DHG-Bench: A Comprehensive Benchmark on Deep Hypergraph Learning

### v1

@@ Abstract
Although conventional deep graph models have achieved great success in relational learning, their focus on pairwise relationships limits their capacity to learn pervasive higher-order interactions in real-world complex systems, which can be naturally modeled as hypergraphs. To tackle this, hypergraph neural networks (HNNs), the dominant approach in deep hypergraph learning (DHGL), has garnered substantial attention in recent years. Despite the proposal of numerous HNN methods, there is no comprehensive benchmark for HNNs, which creates a great obstacle to understanding the progress of DHGL in several aspects: (i) insufficient coverage of datasets, algorithms, and tasks; (ii) a narrow evaluation of algorithm performance; and (iii) inconsistent dataset usage, preprocessing, and experimental setups that hinder comparability. To fill the gap, we introduce DHG-Bench, the first comprehensive benchmark for DHGL. Specifically, DHG-Bench integrates 20 diverse datasets spanning node-, edge-, and graph-level tasks, along with 16 state-of-the-art HNN algorithms, under consistent data processing and experimental protocols. Our benchmark systematically investigates the characteristics of HNNs in terms of four dimensions: effectiveness, efficiency, robustness, and fairness. Further, to facilitate reproducible research, we have developed an easy-to-use library for training and evaluating different HNN methods. Extensive experiments conducted with DHG-Bench reveal both the strengths and inherent limitations of existing algorithms, offering valuable insights and directions for future research. The code is publicly available at: https://github.com/Coco-Hut/DHG-Bench.
@@ 1 Introduction
Graphs are powerful tools for capturing structural and relational properties in various domains, such as social networks (Fan et al., 2019), e-commerce networks (Liu et al., 2021), and financial networks (Li et al., 2024b). Graph Neural Networks (GNNs) have emerged as the dominant approach for learning such data, owing to their strong ability to capture informative topological representations. However, many real-world systems involve multi-way or group-wise interactions beyond the pairwise connections of graphs. For instance, multiple authors co-write a paper in co-authorship networks (Yang et al., 2022), and groups of proteins interact collectively in biological systems (Kim et al., 2024b). These higher-order interactions are naturally modeled by hypergraphs, where each hyperedge connects an arbitrary number of nodes. Directly modeling such higher-order relations has shown improvements across various machine learning tasks compared to graph-based models (Benson et al., 2016; Li & Milenkovic, 2017; Srinivasan et al., 2021; Wang et al., 2023b). This motivates the generalization of GNNs to hypergraphs.
@@
Building on this need, deep hypergraph learning (DHGL) has been developed for learning higher-order knowledge, with hypergraph neural networks (HNNs) emerging as its principal methodology and attracting considerable research interest in recent years. These methods target using neural functions to transform given nodes, hyperedges, and their features into vector representations (i.e., embeddings) that effectively preserve higher-order information. HNNs have shown state-of-the-art performance in industrial and scientific applications, including product recommendation (Ji et al., 2020), 3D object detection (Gao et al., 2012), disease prediction (Gao et al., 2015), and more (Tian et al., 2009; Xu et al., 2018; Zhao et al., 2023).
Despite the emerging studies of HNN algorithms, there lacks a comprehensive and unified benchamrk. HyFER (Hwang et al., 2021) is a modular, easy-to-use framework that supports efficient implementation and fair evaluation of HNNs across diverse datasets, models, and tasks. THU-DH (Gao et al., 2022) is an open-source PyTorch-based toolbox designed for general HNNs. It supports various hypergraph pre-processing methods (e.g., sampling, expansion) and convolution operators, facilitating the evaluation of HNNs. However, existing DHGL benchmarks still suffer from several limitations: (i) The scope of the datasets, algorithms, and benchmark tasks considered in existing studies remains insufficiently comprehensive. For example, HyFER integrates only three HNN methods, while THU-DH supports only approaches published before 2022, thereby excluding many recent state-of-the-art methods. Moreover, the employed datasets are predominantly node-level homophilic hypergraphs, overlooking both heterophilic hypergraphs and those designed for hypergraph-level tasks. This narrow coverage restricts the evaluation of model generalizability across diverse structural patterns. (ii) Algorithm evaluation in existing studies is largely limited to the effectiveness dimension, offering an incomplete picture of model performance and characteristics. Beyond effectiveness, investigating the efficiency and complexity of the algorithm is imperative, yet often overlooked in the literature. Moreover, the trustworthiness of algorithms, particularly their robustness and fairness, remains under-explored despite being crucial for determining their suitability in industrial applications. (iii) The use of different benchmark datasets, data processing approaches, and experimental setups (e.g., parameter settings, initialization procedures) hinders reproducibility and fair comparison across studies. Hence, there is an urgent necessity within the community to develop a standardized and comprehensive benchmark for DHGL.
To bridge this gap, we introduce DHG-Bench, which serves the first open-sourced and comprehensive benchmark for HNNs. First, DHG-Bench implements a wide range of HNN algorithms through unified APIs, while also adopting consistent data processing and data splitting approahces for fair comparisons. Our benchmark encompasses 16 state-of-the-art HNN algorithms and 20 diverse hypergraph datasets covering node-level, edge-level, and graph-level tasks. Moreover, we systematically investigate the characteristics of HNNs in terms of four dimensions: effectiveness, efficiency, robustness, and fairness. Based on these four dimensions, we conduct an extensive comparative study of existing HNN algorithms in different scenarios. For effectiveness, DHG-Bench offers a fair and comprehensive comparison of HNN algorithms on node- and edge-level tasks involving both homophilic and heterophilic hypergraphs. It is also the first benchmark to incorporate the hypergraph-level classification task into evaluation, enabling a more complete assessment of model capabilities across different task levels. For efficiency, DHG-Bench empirically evaluates the efficiency of representative HNNs in terms of both time and space complexity. For robustness, DHG-Bench evaluates HNN models under perturbations at three levels: hypergraph structure, node features, and supervision signals. For fairness, DHG-Bench evaluates the demographic fairness of existing HNNs through a fair node classification task, revealing their performance disparities across different sensitive groups.
Through extensive experiments, we have the following key insights: (1) HNN models show substantial performance variability across different datasets and tasks, and recent approaches have yet to demonstrate consistently better generalization across them. (2) While most HNN methods struggle to strike a satisfactory balance among predictive performance, efficiency, and scalability, TFHNN, by decoupling message passing from model training, achieves the most favorable trade-off. (3) HNN algorithms are susceptible to data perturbations from different aspects, particularly from supervision signals. (4) Although HNNs achieve improved predictive performance, they suffer from significant fairness issues. Our main contributions are summarized as follows:
@@
DHG-Bench enables a fair and unified comparison among 16 state-of-the-art HNN algorithms and 20 diverse hypergraph datasets covering node-level, edge-level, and graph-level tasks. To the best of our knowledge, this is the first comprehensive benchmark for deep hypergraph learning.
@@
We conduct a systematic analysis of HNN methods from various dimensions, including effectiveness, efficiency, robustness, and fairness. Through extensive experiments, we reveal both the strengths and limitations of existing algorithms, offering valuable insights to inform and inspire future research in this field.
@@
We release DHG-Bench, an easy-to-use open-source benchmark library to support future DHGL research. Besides, with our toolkit, users can readily evaluate their algorithms or datasets with less effort.
@@ 2 Preliminary
In this section, we first introduce the notations used throughout the paper, and then give the concepts of three benchmark tasks spanning node-, hyperedge-, and hypergraph-level prediction.
Notations. Let represent a hypergraph with vertex set = and hyperedge set = . denotes the dimensional node feature matrix. As each hyperedge is a subset of , we denote as the size of .
Node Classification. Given the labelded node set and their labels , where each node is associated with a label belonging to one of classes. Semi-supervised node classification aims to train a node classifier to predict labels of the remaining nodes .
Hyperedge Prediction. Given a hypergraph , we deonte as the target set which typically consists of (a) unobserved hyperedges or (b) new hyperedges that will arrive in the near future. Each element in is referred to as a hyperedge candidate, denoted by , as it may belong to . Hyperedge prediction aims to train a hyperedge classifier to preidct whether candidate belongs to the target set or not.
Hypergraph Classification. Let as the hypergraph set. Given the labeled hypergraph set and their labels , where each hypergraph is assigned a label . Semi-supervised hypergaph classification task aims to train a hypergraph classifier to predict labels of the unlabeled hypergraphs .
@@ 3 Benchmark Design
In this section, we introduce the DHG-Bench in terms of datasets (Section 3.1), algorithms (Section 3.2), and research questions (Section 3.3) that guide the benchmark design. The overview of DHG-Bench is illustrated in Figure 1.
@@ 3.1 Benchmark Datasets
To comprehensively and effectively evaluate the characteristics of various HNN algorithms, we have integrated 20 real-world datasets from various domains for node-level, hyperedge-level, and hypergraph-level tasks. We briefly introduce each category in this section and provide detailed descriptions in Appendix A.1.
@@
Node-level Classification Datasets. For the most prominent task in hypergraph learning, node classification, we utilize 11 hypergraph datasets spanning diverse domains and exhibiting varying characteristics. As summarized in Table A1 in Appendix A, the benchmark includes 6 homophilic datasets: 2 co-citation networks (Cora, Pubmed (Yadati et al., 2019)), 2 co-authorship networks (Cora-CA, DBLP (Yadati et al., 2019)), and 2 large-scale benchmarks (Walmart (Chien et al., 2022), and Trivago (Kim et al., 2023)) that model co-purchasing, and co-click interactions, respectively. In addition, we incorporate 5 heterophilic datasets that capture multi-way relationships among entities in various contexts, including two information networks (Actor (Li et al., 2025) and Yelp (Chien et al., 2022)), an e-commerce network (Amazon-ratings (Li et al., 2025)), and two social networks (Twitch-gamers, Pokec (Li et al., 2025)). To further investigate potential fairness concerns in existing HNN algorithms, we include three fairness-sensitive datasets (German, Bail, and Credit (Wei et al., 2022)), which contain sensitive node attributes such as gender, race, and age. Detailed statistics of these datasets are presented in Table A2. In line with previous studies (Feng et al., 2019; Chien et al., 2022; Wang et al., 2023a; Tang et al., 2025), we adopt a 50%/25%/25% split for training, validation, and testing in the node classification task.
@@
Hyperedge-level Prediction Datasets. For the hypergraph prediction task, we consider 4 widely used real-world datasets: two co-citation networks (Cora, Pubmed) and two co-authorship networks (Cora-CA, DBLP-CA), as adopted in prior studies (Hwang et al., 2022; Ko et al., 2025). To support a more comprehensive evaluation, we further include two heterophilic datasets, Actor and Pokec, characterized by low hyperedge homophily (Li et al., 2025). Following the protocol in (Hwang et al., 2022; Ko et al., 2025; Yu et al., 2025), we randomly split the hyperedges (i.e., positive samples) into training (60%), validation (20%), and test (20%) sets, repeated over five random seeds. To rigorously assess the performance of various HNNs, we construct four different validation and test sets, each of which has different negative examples with various degrees of difficulty. Specifically, for each validation and test set, we sample an equal number of negative examples as the positive ones. Following [1], we adopt a mixed negative sampling (NS) strategy that incorporates three widely adopted heuristic NS methods: sized NS (SNS), motif NS (MNS), and clique NS (CNS) (Patil et al., 2020), thereby increasing the diversity of negative samples.
@@
Hypergraph-level Classification Datasets. For the hypergraph classification task, We consider 6 benchmark datasets introduced in a recent study (Feng et al., 2024), as summarized in Table A3. RHG-10 is a synthetic datasets which encompass ten distinct hypergraph structure factors (e.g., Hyper Flower and Hyper Pyramid). RHG-3 is a synthetic dataset consisting of randomly generated hypergraphs with three distinctively various hypergraph structures: hyper-pyramid, hyper-check-table, and hyper-wheel. IMDB-Dir-Form and IMDB-Dir-Genre are two datasets constructed by the co-director relationship from the original IMDB dataset 111https://www.imdb.com/. Steam-Player is a player-based dataset, where each hypergraph captures tag co-occurrence relationships among games played by a user. Twitter-Friend is a social media dataset, where each hypergraph represents the friendship network of a specific Twitter user. For hypergraph classification, we adopt a 80%/10%/10% train/validation/test data split.
@@ 3.2 Benchmark Algorithms
We have integrated 16 representative and competitive HNN algorithms in DHG-Bench. They are divided into three technique categories: Spectral-based HNN, Spatial-based HNN, and Tensor-based HNN. We briefly introduce each category in the following, and more details are provided in Appendix A.2.
Spectral-based HNN Algorithms. Spectral-based HNN methods are a class of algorithms that leverage the spectral theory of hypergraph Laplacians to perform message propagation and feature transformation. These models operate in the spectral domain, where convolution operations are defined based on the eigendecomposition of a hypergraph Laplacian matrix derived from the adjacency matrix of the hypergraph (Saxena et al., 2024; Wang et al., 2024a). We integrate 10 representative spectral-based algorithms including HGNN (Feng et al., 2019), HyperGCN (Yadati et al., 2019), HCHA (Bai et al., 2021), LEGCN (Yang et al., 2022), HyperND (Prokopchik et al., 2022), PhenomNN (Wang et al., 2023b), SheafHyperGNN (Duta et al., 2023), HJRL (Yan et al., 2024), DPHGNN (Saxena et al., 2024), and TF-HNN (Tang et al., 2025).
Spatial-based HNN Algorithms. In contrast to spectral-based method, spatial-based methods focus on the local connectivity of each node without going to the spectral domain (Wang et al., 2024a). They commonly apply a two-stage neighborhood aggregation scheme to learn hypergraph representations, which entails updating the hyperedge embedding by aggregating embeddings of its incident nodes, and updating the node representation by propagating information from representations of its incident hyperedges (Gao et al., 2022; Wang et al., 2024a). We consider 4 spatial-based algorithms including HNHN (Dong et al., 2020), UniGNN (Huang & Yang, 2021), AllSetTransformer (Chien et al., 2022), and ED-HNN (Wang et al., 2023a). Since multiple UniGNN variants exist (e.g., UniGAT, UniGIN, UniGCNII), we report only the results of the most competitive variant, UniGCNII, for simplicity. We also note that our open-source library supports the implementations of other variants.
Tensor-based HNN Algorithms. These methods leverage tensor operations that provide a structured and effective means of capturing the complexity of hypergraph interactions (Wang et al., 2024b; 2025). In our DHG-Bench, we integrate two representative tensor-based algorithms: EHNN (Kim et al., 2022), and T-HyperGNN (Wang et al., 2024a).

### v2

# DHG-Bench: A Comprehensive Benchmark for Deep Hypergraph Learning

@@ Abstract
Deep graph models have achieved great success in network representation learning. However, their focus on pairwise relationships restricts their ability to learn pervasive higher-order interactions in real-world systems, which can be naturally modeled as hypergraphs. To tackle this issue, Hypergraph Neural Networks (HNNs) have garnered substantial attention in recent years. Despite the proposal of numerous HNNs, the absence of consistent experimental protocols and multi-dimensional empirical analysis impedes deeper understanding and further development of HNN research. While several toolkits for deep hypergraph learning (DHGL) have been introduced to facilitate algorithm evaluation, they provide only limited quantitative evaluation results and insufficient coverage of advanced algorithms, datasets, and benchmark tasks. To fill the gap, we introduce DHG-Bench, the first comprehensive benchmark for HNNs. Specifically, DHG-Bench systematically investigates the characteristics of HNNs in terms of four dimensions: effectiveness, efficiency, robustness, and fairness. We comprehensively evaluate 17 state-of-the-art HNN algorithms on 22 diverse datasets spanning node-, edge-, and graph-level tasks, under unified experimental settings. Extensive experiments reveal both the strengths and limitations of existing algorithms, offering valuable insights and directions for future research. Furthermore, to facilitate reproducible research, we have developed an easy-to-use library for training and evaluating different HNN methods. The DHG-Bench library is available at: https://github.com/Coco-Hut/DHG-Bench.
@@ 1 Introduction
Graph-structured data has become a ubiquitous tool for modeling the complex relational dependencies among entities in various domains, such as social analysis (Fan et al., 2019), e-commerce (Liu et al., 2021), and finance (Li et al., 2024). Graph Neural Networks (GNNs) have emerged as the dominant approach for learning on such data, owing to their exceptional ability to leverage both the graph topology and node attributes. However, many real-world systems involve multi-way or group-wise interactions beyond the pairwise connections of graphs. For instance, multiple authors co-write a paper in co-authorship networks (Yang et al., 2022), and groups of proteins interact collectively in biological systems (Kim et al., 2024). These higher-order interactions can be naturally modeled by hypergraphs, where each hyperedge connects an arbitrary number of nodes. As hypergraphs become increasingly prevalent, there is a growing demand for predictive tasks on them, such as estimating node properties or identifying missing hyperedges (Kim et al., 2024). However, directly applying GNNs to such tasks inevitably collapses higher-order interactions into pairwise relations, resulting in significant information loss and thus sub-optimal performance (Chien et al., 2022).
@@
To mitigate the aforementioned problem, Hypergraph Neural Networks (HNNs) (Yadati et al., 2019; Chien et al., 2022; Wang et al., 2023b; Tang et al., 2025) have become the prevailing paradigm for deep hypergraph learning (DHGL), attracting considerable research interest in recent years. These methods employ neural architectures to transform nodes, hyperedges, and their associated features into vector representations (i.e., embeddings) that effectively preserve higher-order semantics. HNNs have demonstrated state-of-the-art performance across diverse industrial and scientific applications, including product recommendation (Khan et al., 2025), 3D object detection (Fixelle, 2025), and disease diagnosis (Han et al., 2025).
Despite the emerging studies of HNN algorithms, the comprehensive benchmark for evaluating these methods remains absent, bringing out the following problems: (i) Existing works utilize different datasets, compared baselines, and experimental setups (e.g., data splitting strategies and parameter settings), which makes it challenging to achieve a fair comparison. (ii) Existing works primarily focus on the effectiveness evaluation of HNN algorithms, while lacking empirical understanding of their efficiency and trustworthiness (e.g., robustness and fairness), both of which are essential for real-world deployment. This prevents practitioners from understanding the advantages and limitations of HNN algorithms from multiple perspectives and makes it difficult to select appropriate methods for different application scenarios. Hence, there is an urgent necessity within the community to develop a standardized and comprehensive benchmark for HNNs.
In recent years, several open-sourced toolkits, including HyFER (Hwang et al., 2021), DHG (Gao et al., 2022), and TopoX (Hajij et al., 2024), have been proposed to facilitate benchmarkable deep hypergraph learning. However, these works provide only limited or even no quantitative performance comparisons, which thus compromises their practical value for practitioners. Furthermore, they fail to incorporate many state-of-the-art HNN algorithms and provide insufficient coverage of benchmark datasets and evaluation tasks. Specifically, HyFER supports only the implementation of three HNN models, while the other two libraries include only HNNs proposed before 2023. Moreover, none of these toolkits integrate heterophilic hypergraph datasets, which represent a particularly challenging setting (Li et al., 2025c), nor do they support graph-level tasks (e.g., hypergraph classification). These limitations significantly restrict the reproducibility and comprehensive evaluation of advanced HNNs.
To bridge the gap, we propose DHG-Bench, which serves as the first open-sourced and comprehensive benchmark for HNNs. Our benchmark encompasses 17 representative HNN methods and 22 diverse hypergraph datasets covering node-level, edge-level, and graph-level tasks. We employ standardized computational operators and APIs, along with consistent data splitting and processing strategies, to ensure fair comparison. Beyond effectiveness, our benchmark supports multi-faceted analysis, allowing researchers to investigate the efficiency, robustness, and fairness of current HNN algorithms. Through extensive experiments, we derive the following key insights: (i) Existing HNN algorithms exhibit substantial performance variability across datasets and tasks, reflecting their limited generalization ability. (ii) Most HNN methods struggle to strike a satisfactory balance between predictive performance and computational efficiency. (iii) The performance of HNN algorithms is affected by different types of data perturbations, with feature-level and supervision-level perturbations causing particularly adverse impacts. (iv) HNN algorithms tend to result in more severe fairness issues than deep models without higher-order message passing, such as MLPs. Our main contributions are summarized as follows:
@@
Comprehensive Benchmark. DHG-Bench enables a fair and unified comparison among 17 state-of-the-art HNN methods by standardizing the experimental settings across 22 widely used hypergraph datasets of diverse characteristics. To the best of our knowledge, this is the first comprehensive benchmark for deep hypergraph learning.
@@
Multi-dimensional Evaluation and Analysis. We conduct a systematic analysis of existing HNN methods from various dimensions, encompassing effectiveness, efficiency, robustness, and fairness. Extensive experiments uncover the potential strengths and limitations of existing HNN algorithms, offering valuable insights to inform and inspire future research in this field.
@@
Open-sourced Benchmark Library. We release DHG-Bench, an easy-to-use open-sourced benchmark library to support future HNN research. With our toolkit, users can evaluate their algorithms or datasets with less effort.
@@ 2 Preliminary
Let represent a hypergraph with vertex set = and hyperedge set = . is the node feature matrix with -dimension. In this benchmark, we focus on three supervised learning tasks, covering node-, edge-, and graph-level prediction.
Node Classification. Given the labeled node set with labels , where each node is associated with a label from one of the classes, the goal of node classification is to train a classifier to predict labels of the remaining unlabeled nodes .
Hyperedge Prediction. Given a hypergraph , we denote as the target set which typically consists of (a) unobserved hyperedges or (b) new hyperedges that will arrive in the near future. Each element in is referred to as a hyperedge candidate, denoted by , as it may belong to . The hyperedge prediction task aims to train a hyperedge classifier to predict whether a candidate belongs to the target set or not.
Hypergraph Classification. Let as the hypergraph set. Given the labeled hypergraph set and their labels , where each hypergraph is assigned a label . The hypergraph classification task aims to train a hypergraph classifier to predict labels of the unlabeled hypergraphs .
@@ 3 Benchmark Design
In this section, we introduce the DHG-Bench in terms of datasets (Section 3.1), algorithms (Section 3.2), and research questions (Section 3.3) that guide the benchmark study.
@@ 3.1 Benchmark Datasets
To comprehensively evaluate HNNs, we integrate 22 benchmark datasets from various domains spanning node-, edge-, and graph-level tasks. In this section, we introduce each dataset category and the corresponding data splitting strategy. Detailed descriptions are provided in Appendix A.1.
@@
Node-level Classification Datasets. For the node classification task, we select 13 hypergraph datasets that cover diverse domains and characteristics. Specifically, we include 8 homophilic datasets: two co-citation networks (Cora and Pubmed (Yadati et al., 2019)); two co-authorship networks (Cora-CA and DBLP (Yadati et al., 2019)); two graphics datasets (NTU2012 and ModelNet40 (Feng et al., 2019)); and two hypergraphs that capture user interactions, namely Walmart for co-purchasing (Chien et al., 2022) and Trivago for co-clicking (Kim et al., 2023). In addition, we consider 5 heterophilic datasets, including two information networks (Actor (Li et al., 2025c) and Yelp (Chien et al., 2022)), an e-commerce network (Amazon-ratings (Li et al., 2025c)), and two social networks (Twitch-gamers and Pokec (Li et al., 2025c)). Moreover, to investigate algorithmic fairness, we include three fairness-sensitive datasets (German, Bail, and Credit (Wei et al., 2022)), which contain sensitive node attributes such as gender, race, and age. Following (Feng et al., 2019; Chien et al., 2022; Tang et al., 2025), we adopt a split of 50%/25%/25% for training, validation, and testing in the node classification task.
@@
Hyperedge-level Prediction Datasets. For the hyperedge prediction task, we use 6 datasets: four widely adopted homophilic academic networks (Cora, Pubmed, Cora-CA, and DBLP-CA) (Hwang et al., 2022; Ko et al., 2025) and two newly introduced heterophilic datasets, Actor and Pokec (Li et al., 2025c), which enable a more comprehensive evaluation due to their low hyperedge homophily. Following (Hwang et al., 2022; Ko et al., 2025; Yu et al., 2025), we randomly split the hyperedges (i.e., positive samples) into training (60%), validation (20%), and test (20%) sets. In addition, we adopt negative sampling (NS) (Yadati et al., 2020; Hwang et al., 2022), which is devised to enhance the distinguishing ability of the model by introducing non-existing hyperedges as contrastive information for model training. Specifically, for each training, validation, and test set, we sample an equal number of negative examples as the positive ones. Following (Ko et al., 2025), we employ a mixed NS strategy that integrates three common heuristic methods, namely sized NS (SNS), motif NS (MNS), and clique NS (CNS) (Patil et al., 2020), to increase the diversity of negative samples.
@@
Hypergraph-level Classification Datasets. For the hypergraph classification task, we consider 6 benchmark datasets introduced in (Feng et al., 2024). RHG-10 and RHG-3 are two synthetic datasets consisting of distinct high-order structural patterns (e.g., Hyper Pyramid, Hyper Flower, and Hyper Wheel). IMDB-Dir-Form and IMDB-Dir-Genre are two datasets constructed by the co-director relationship from the original IMDB dataset 111https://www.imdb.com/. Steam-Player is a player-based dataset, where each hypergraph captures tag co-occurrence relationships among games played by a user. Twitter-Friend is a social media dataset where each hypergraph represents the friendship network of a specific Twitter user. For hypergraph classification, following (Feng et al., 2024), we adopt an 80%/10%/10% train/validation/test data split.
@@ 3.2 Benchmark Algorithms
We integrate 17 state-of-the-art HNN algorithms across three mainstream categories: spectral-based, spatial-based, and tensor-based methods. In addition, we include MLP and two GNN-based methods, CEGCN and CEGAT (Chien et al., 2022), as baselines. Detailed descriptions are provided in Appendix A.2. We rigorously reproduce all methods according to their papers and source codes.
Spectral-based HNNs. Spectral-based HNNs perform message propagation and feature transformation by applying spectral convolution defined through Laplacian operators of hypergraphs (Wang et al., 2024). We implement 10 representative algorithms including HGNN (Feng et al., 2019), HyperGCN (Yadati et al., 2019), HCHA (Bai et al., 2021), LEGCN (Yang et al., 2022), HyperND (Prokopchik et al., 2022), PhenomNN (Wang et al., 2023b), SheafHyperGNN (Duta et al., 2023), HJRL (Yan et al., 2024), DPHGNN (Saxena et al., 2024), and TF-HNN (Tang et al., 2025).
Spatial-based HNNs. Unlike spectral methods, spatial-based HNNs focus on local connectivity without entering the spectral domain, typically learning representations through two-stage neighborhood aggregation: updating hyperedges from incident nodes and updating nodes from incident hyperedges. We incorporate 5 typical algorithms including HNHN (Dong et al., 2020), UniGNN (Huang & Yang, 2021), AllSetTransformer (Chien et al., 2022), ED-HNN (Wang et al., 2023a), and HyperGT (Liu et al., 2024). For UniGNN with multiple variants (e.g., UniGAT, UniGIN, and UniGCNII), we report only UniGCNII, the most competitive variant identified in the original paper, while our open-sourced library also supports the implementations of other variants.